OpenStudy (anonymous):

Simplify. Express the quotient as a radical expression. ^9sqrt x^7 / ^9sqrt x^5 Simplify. Express the quotient as a radical expression. ^9sqrt x^7 / ^9sqrt x^5 @Mathematics

OpenStudy (agreene):

\[\large \frac{\sqrt[9] {x^7}}{\sqrt[9] {x^5}}\] Is this the question?

OpenStudy (anonymous):

yeah

OpenStudy (agreene):

Hrm... The only thing I can think of to reduce this is: \[\large\forall x\ge0\exists f(x)|x^{\frac{2}{9}}\]

OpenStudy (anonymous):

i can give you my answers if that'll help.

OpenStudy (agreene):

Oh it just wants the quotient rationalized... \[\large\frac{(x^5)^{8/9}\sqrt[5]{x^7}}{x^5}\]

OpenStudy (agreene):

By quotient I meant denominator... and you can just pull 8/9ths of the 1/9th total and pull to the top, those are equivocalness.

OpenStudy (agreene):

Wow, I cant type tonight--I think I should go to sleep soon O_o

OpenStudy (agreene):

But yeah, if you have some possible answers, shoot 'em at me and I'll see which one is the closest to those two answers.

OpenStudy (anonymous):

\[\sqrt[9]{x ^{2}}\] \[\sqrt[0]{x ^{2}}\] \[\sqrt[9]{x ^{12}}\] \[\sqrt[81]{x ^{35}}\]

OpenStudy (agreene):

If one of those is 9th root of x^(-2) then that kind of makes sense, if you forget how roots actually work. This is kind of annoying because none of these are true... if there isnt a -2 power, then id go with the 9^sqrt x^2 I'd point out to your professor that this question is not a true statement.

OpenStudy (anonymous):

okay, thank you sososososo much.